# Near Optimal Non-asymptotic Sample Complexity of 1-Identification

### v1

@@ 1 Introduction
Notation. For an integer $K>0$, denote $[K]=\{1,\ldots,K\}$. For $\mu\in[0,1]$, we denote $N(\mu,\sigma^{2})$ as the normal distribution with mean $\mu$ and variance $\sigma^{2}$. Denote $\mathbb{E}$ and $\mathbb{E}_{\nu,\textsf{alg}}$ as the expectation operator while the second one is to highlight that the expectation is determined by both the instance $\nu$ and algorithm alg.
@@ 2 Model
Dynamics. The agent’s pulling strategy $(\pi,\tau,\hat{a})$ is parametrized by a sampling rule $\pi=\{\pi_{t}\}_{t=1}^{\infty}$, a stopping time $\tau$ and a recommendation rule $\hat{a}\in[K]\cup\{\textsf{None}\}$. When a pulling strategy is applied on a 1-identification problem instance $\nu$, the strategy generates a history $A_{1},X_{1},\cdots,A_{\tau},X_{\tau}$. Action $A_{t}\in[K]$ is chosen contingent upon the history $H(t-1)=\{(A_{s},X_{s})\}^{t-1}_{s=1}$, via the function $\pi_{t}(H(t-1))$. In addition, we have $X_{t}\sim\nu_{A_{t}}$. The agent stops pulling at the end of time step $\tau$, where $\tau$ is a stopping time [^1] For any $t$, the event $\{\tau=t\}$ is $\sigma(H(t))$-measurable with respect to the filtration $\{\sigma(H(t))\}^{\infty}_{t=1}$. Upon stopping, the agent outputs $\hat{a}\in[K]\cup\{\textsf{None}\}$ to be the answer, using the information $H(\tau)$. Outputting $\hat{a}\in[K]$ means that the agent concludes with arm $\hat{a}$ satisfying $\mu_{\hat{a}}\geq\mu_{0}$, while outputting $\hat{a}=\textsf{None}$ means that the agent concludes with no arm having mean reward $\geq\mu_{0}$. We allow the possibility of non-termination $\tau=\infty$, in which case there is no recommendation $\hat{a}$.
@@
Objective. The agent aims to design a $(\Delta,\delta)$-PAC pulling strategy $(\pi,\tau,\hat{a})$ that minimizes the sampling complexity $\mathbb{E}[\tau]$.
@@ 4.1 An Informal Algorithm
5:     • Total pulling times in all exploration phases is not greater than $T_{k}^{\text{ee}}$. Take $\hat{a}_{k}=\textsf{Not Complete}$
@@
9:        • Pulling times of $\hat{a}_{k}$ is not smaller than $T_{k}^{\text{et}}$.
@@
Algorithm 1 introduces the key idea of SEE. Algorithm LUCB_G in line 4 is defined in (Kano et al., 2017), which can be considered an adapted BAI algorithm UCB in (Jamieson & Nowak, 2014). In each round, LUCB_G pulls the arm with the highest upper confidence bound (UCB), and stops if there is an arm whose lower confidence bound (LCB)is $>\mu_{0}$. LUCB_G guarantees the high probability bound $\Pr_{\nu}(\mu_{\hat{a}}>\mu_{0},\tau<O(H_{1}^{\text{pos}}\log(H_{1}^{\text{pos}}/\delta^{\prime})))>1-\delta^{\prime}$ for $\nu\in\mathcal{S}^{\text{pos}}$, and $\Pr_{\nu}(\hat{a}=\textsf{None},\tau<O(H_{1}^{\text{neg}}\log(H_{1}^{\text{neg}}/\delta^{\prime})))>1-\delta^{\prime}$ for $\nu\in\mathcal{S}^{\text{neg}}$, where $\delta^{\prime}\in(0,1)$ is the input tolerance level.
@@
The second concern is LUCB_G requires at the start of an exploration phase $k$, $\text{LCB}_{a}(\delta_{k})<\mu_{0}$ holds for all $a\in[K]$. But in Algorithm 1, it is possible that at the end of phase $k-1$, the last collected sample of arm $\hat{a}_{k-1}$ is so large, such that LCB of $\hat{a}_{k-1}$ is still above $\mu_{0}$ at the start of phase $k$. To address this issue, we define a temporary container $Q$. When the LUCB_G is going to output an arm $\hat{a}_{k-1}$, we transfer the latest collected sample of $\hat{a}_{k-1}$ into $Q$. If we pull arm $\hat{a}_{k-1}\in[K]$ in the future exploration period, we transfer the sample back from $Q$ to history. The reason of transferring back from $Q$ is concentration inequality like (3) requires consecutive integer index summation, and we cannot skip or abandon any samples.
@@ 4.2 Sequential Exploration Exploitation
Algorithm 4 takes $(K,\mu_{0},\delta,T,\hat{a})$ as the input. The parameters $K,\mu_{0},\delta$ are model parameters, which remain the same in all phases. By contrast, the parameters $T$ and $\hat{a}$ generally change across different phases. During phase $k$, Algorithm 4 keeps pulling arm $\hat{a}_{k}$ until one of the two following conditions is met. The first is when the total pulling times of arm $\hat{a}_{k}$ in all exploitation periods is $T$, which should be $T_{k}^{\text{et}}$ at the phase $k$. In this case, the exploitation period cannot guarantee that $\mu_{\hat{a}_{k}}>\mu_{0}$ holds with probability $\geq 1-\delta$, and requires Algorithm 2 to run the next exploration period with a smaller tolerance level $\delta_{k+1}$. The second condition is when the LCB, defined with the input tolerance level $\delta$, of $\hat{a}_{k}$ is above $\mu_{0}$. In this case, Algorithm 2 adopts this result and outputs $\hat{a}_{k}$ as $\hat{a}$, asserting that the instance is positive.
@@ 5.1 Performance Guarantee of Algorithm 2
If $\nu$ is positive, the event $\hat{a}\notin i^{*}(\nu)$ occurs only in the following two scenarios. The first scenario is when an exploration period outputs None in phase $k$ such that $\delta_{k}<\delta/3$. The second scenario is when an exploitation period outputs an arm with mean reward $<\mu_{0}$. The probabilities of both events are at most $\delta$ multiplied by an absolute constant, and we can conclude $\Pr_{\nu}(\hat{a}\notin i^{*}(\nu))<\delta$.
@@ 6 Numerical Experiments
We consider six different groups of reward vectors, which are “All Worse”, “Unique Qualified”, “One Quarter”, “Half Good”, “All Good” and “Linear”. The main difference among these groups is the number of qualified arms. All instances in “All Worse” group are negative, whose mean rewards of all arms are below $\mu_{0}$. Due to the low execution speed of algorithm MS on group “All Worse”, we do not apply MS to the instance “All Worse”. Other groups of instances are positive, meaning that there is at least one arm whose mean reward is $>\mu_{0}$. We set $10^{8}$ as our forced stopping threshold for each group of experiments. Every algorithm stops and outputs an answer at or before the number of arm pulls reaches $10^{8}$. The details of the numerical setting, tuning of the hyper-parameters in SEE, and other implementation details can be found in Appendix E.1.
@@ B.2 Proof of Main Theorems
$O\left(K\beta_{\tilde{L}^{\prime}}\log\frac{4K}{\delta_{\tilde{L}^{\prime}}}\right)=$ $O\left(H_{1}^{\text{neg}}\log H_{1}^{\text{neg}}\right)$
@@
$O(H_{1}^{\text{neg}}\log\frac{1}{\delta_{\tilde{L}^{\prime}}})\leq$ $O\left(H_{1}^{\text{neg}}\log H_{1}^{\text{neg}}\right)$
$O(H_{1}^{\text{neg}}\log\frac{1}{\delta_{\tilde{L}^{\prime\prime}}})\leq$ $O\left(H_{1}^{\text{neg}}\log\frac{1}{\delta}\right)$
@@ B.3 Supplement Lemmas
From the algorithm design, we know the total pulling times in the exploration period up to phase $k-1$ is at most $T_{k-1}^{\text{ee}}$. By the Lemma B.6, we can validate the conditions in Lemma B.4 holds for $k\geq\max\left\{\kappa^{\text{ee}},\lceil\log_{2}\frac{24(C+1)^{2}H_{1}^{\text{pos}}}{K}\rceil\right\}$. Thus, we can assert $\hat{a}_{k}\in[K],\mu_{\hat{a}_{k}}\geq\omega\mu_{1}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$ for $k\geq L$. The remaining work is to prove the upper bound of $\tau_{k}^{\text{ee}}$.
@@
From the algorithm design, we know the total pulling times in the exploration period up to phase $k-1$ is at most $T_{k-1}^{\text{ee}}$. By the Lemma B.5, we can validate the conditions of Lemma B.5 hold for $k\geq\left\{\kappa^{\text{ee}},\lceil\log_{2}\frac{\sum_{a=1}^{K}\frac{24}{\Delta_{0,a}^{2}}}{K}\rceil\right\}$. Thus, we have proved $\hat{a}_{k}=\textsf{None}$ for $k\geq\left\{\kappa^{\text{ee}},\lceil\log_{2}\frac{\sum_{a=1}^{K}\frac{24}{\Delta_{0,a}^{2}}}{K}\rceil\right\}$. The remaining work is to prove the upper bound of $N_{a}(\tau_{k}^{\text{ee}})$.
@@ C.3 Lower Bound for a Suboptimal Arm
where $Z_{1,s}$ is the realized KL divergence, $Z_{1,s}=\log\frac{\exp(-\frac{(X_{1,s}-\mu_{1})^{2}}{2})}{\exp(-\frac{(X_{1,s}-\tilde{\mu}_{1})^{2}}{2})}=\frac{(\mu_{1}-\tilde{\mu}_{1})(2X_{1,s}-\mu_{1}-\tilde{\mu}_{1})}{2}$, $X_{1,s}\sim N(\mu_{1},1)$. Or we can directly assume $Z_{1,s}\stackrel{{\scriptstyle i.i.d}}{{\sim}}N(\frac{(\mu_{1}-\tilde{\mu}_{1})^{2}}{2},(\mu_{1}-\tilde{\mu}_{1})^{2})$. Before moving on, we firstly prove a concentration result for the sum $\sum_{s=1}^{t}Z_{1,s}$. For any fixed number $N$, $I$, define $\xi_{N,I}=\left\{\max_{1\leq t\leq N}\sum_{s=1}^{t}\left(Z_{1,s}-\frac{(\mu_{1}-\tilde{\mu}_{1})^{2}}{2}\right)<I\right\}$. We can assert $\Pr_{Z_{1,s}\stackrel{{\scriptstyle i.i.d}}{{\sim}}N(\frac{(\mu_{1}-\tilde{\mu}_{1})^{2}}{2},(\mu_{1}-\tilde{\mu}_{1})^{2})}(\xi_{N,I})\geq 1-\exp\left(-\frac{I^{2}}{2N(\mu_{1}-\tilde{\mu}_{1})^{2}}\right)$, by the following application of maximal inequality (Theorem 3.10 in (Lattimore & Szepesvári, 2020)).
@@
=$\stackrel{{\scriptstyle\lambda=\frac{I}{N(\mu_{1}-\tilde{\mu}_{1})^{2}}}}{{=}}$ $\Pr_{\tilde{Z}_{1,s}\stackrel{{\scriptstyle i.i.d}}{{\sim}}N(0,(\mu_{1}-\tilde{\mu}_{1})^{2})}\left(\max_{1\leq t\leq N}\prod_{s=1}^{t}\exp\left(\lambda\tilde{Z}_{1,s}\right)>\exp(\lambda I)\right)$
@@ D.1 Inequality about $x$ and $\log\log x$
>$\stackrel{{\scriptstyle x>2\log x,\forall x>0}}{{>}}$ $0,$
@@ D.2 Probability Bound of Good Event
=$\stackrel{{\scriptstyle\lambda=\frac{x}{u_{k}\sigma^{2}}}}{{=}}$ $\exp(-\frac{x^{2}}{2u_{k}\sigma^{2}}).$
@@ E.1 Settings of Numerical Experiments
[…]
@@ E.3 Correctness of APGAI algorithm
Instance Type $\setminus$ K $10$ $20$ $30$ $40$ $50$
@@
Instance Type $\setminus$ K $10$ $20$ $30$ $40$ $50$
@@
We tune some of the numeric experiment setting to see the non-stopping phenomenon more clearly. For APGAI, we set the forced stopping threshold as $50000*K$, where $K$ is the arm number. Once the total pulling times is no less than the forced stopping threshold, we mark this experiment as a failure, and take the forced stopping threshold as the total pulling times of APGAI. Table 2 records the number of failure for experiments $K=10,20,30,40,50$ and instance type “All Worse”, “Unique Qualified”, “One Quarter”, “Linear”. The tolerance level is $\delta=0.001$, with repeating times 1000. Except the negative instance “All Worse”, failure frequently occurs in some of positive instances. In the “Unique Qualified” group, at least 35% of experiments end up with “failure”. While in other groups, failure also occurs. In group “One Quarter”, the number of failures decreases as arm number $K$ increases, suggesting APGAI’s good performance relies more on the number of qualified arms, instead of the fraction of qualified arms.

### v2

@@ 1 Introduction
Notation. For an integer $K>0$, denote $K]=\{1,\ldots,K\}$. For $\mu\in[0,1]$, we denote $N(\mu,\sigma^{2})$ as the normal distribution with mean $\mu$ and variance $\sigma^{2}$. Denote $\mathbb{E}$ and $\mathbb{E}_{\nu,\textsf{alg}}$ as the expectation operator while the second one is to highlight that the expectation is determined by both the instance $\nu$ and algorithm alg.
@@ 2 Model
[…]
@@
[…]
@@ 4.1 An Informal Algorithm
5:  • Total pulling times in all exploration phases is greater than $T_{k}^{\text{ee}}$. Take $\hat{a}_{k}=\textsf{Not Complete}$
@@
9:   • Pulling times of $\hat{a}_{k}$ is greater than $T_{k}^{\text{et}}$.
@@
Algorithm 1 introduces the key idea of SEE. Algorithm LUCB_G in line 4 is defined in (Kano et al., 2017), which can be considered an adapted BAI algorithm UCB in (Jamieson & Nowak, 2014). In each round, LUCB_G pulls the arm with the highest upper confidence bound (UCB), and stops if there is an arm whose lower confidence bound (LCB)is $\mu_{0}$. LUCB_G guarantees the high probability bound $\Pr_{\nu}(\mu_{\hat{a}}>\mu_{0},\tau<O(H_{1}^{\text{pos}}\log(H_{1}^{\text{pos}}/\delta^{\prime})))>1-\delta^{\prime}$ for $\nu\in\mathcal{S}^{\text{pos}}$, and $\Pr_{\nu}(\hat{a}=\textsf{None},\tau<O(H_{1}^{\text{neg}}\log(H_{1}^{\text{neg}}/\delta^{\prime})))>1-\delta^{\prime}$ for $\nu\in\mathcal{S}^{\text{neg}}$, where $\delta^{\prime}\in(0,1)$ is the input tolerance level.
@@
The second concern is LUCB_G requires at the start of an exploration phase $k$, $\text{LCB}_{a}(\delta_{k})<\mu_{0}$ holds for all $a\in[K]$. But in Algorithm 1, it is possible that at the end of phase k1k-1, the last collected sample of arm $\hat{a}_{k-1}$ is so large, such that LCB of $\hat{a}_{k-1}$ is still above $\mu_{0}$ at the start of phase $k$. To address this issue, we define a temporary container $Q$. When the LUCB_G is going to output an arm $\hat{a}_{k-1}$, we transfer the latest collected sample of $\hat{a}_{k-1}$ into $Q$. If we pull arm $\hat{a}_{k-1}\in[K]$ in the future exploration period, we transfer the sample back from $Q$ to history. The reason of transferring back from $Q$ is concentration inequality like (3) requires consecutive integer index summation, and we cannot skip or abandon any samples.
@@ 4.2 Sequential Exploration Exploitation
[…]
@@ 5.1 Performance Guarantee of Algorithm 2
If $\nu$ is positive, the event $\hat{a}\notin i^{*}(\nu)$ occurs only in the following two scenarios. The first scenario is when an exploration period outputs None in phase $k$ such that $\delta_{k}<\delta/3$. The second scenario is when an exploitation period outputs an arm with mean reward $\mu_{0}$. The probabilities of both events are at most $\delta$ multiplied by an absolute constant, and we can conclude $\Pr_{\nu}(\hat{a}\notin i^{*}(\nu))<\delta$.
@@ 6 Numerical Experiments
We consider six different groups of reward vectors, which are “All Worse”, “Unique Qualified”, “One Quarter”, “Half Good”, “All Good” and “Linear”. The main difference among these groups is the number of qualified arms. All instances in “All Worse” group are negative, whose mean rewards of all arms are below $\mu_{0}$. Due to the low execution speed of algorithm MS on group “All Worse”, we do not apply MS to the instance “All Worse”. Other groups of instances are positive, meaning that there is at least one arm whose mean reward is $\mu_{0}$. We set $10^{8}$ as our forced stopping threshold for each group of experiments. Every algorithm stops and outputs an answer at or before the number of arm pulls reaches $10^{8}$. The details of the numerical setting, tuning of the hyper-parameters in SEE, and other implementation details can be found in Appendix E.1.
@@ B.2 Proof of Main Theorems
[…]
@@
[…]
@@ B.3 Supplement Lemmas
From the algorithm design, we know the total pulling times in the exploration period up to phase k1k-1 is at most $T_{k-1}^{\text{ee}}$. By the Lemma B.6, we can validate the conditions in Lemma B.4 holds for $k\geq\max\left\{\kappa^{\text{ee}},\lceil\log_{2}\frac{24(C+1)^{2}H_{1}^{\text{pos}}}{K}\rceil\right\}$. Thus, we can assert $\hat{a}_{k}\in[K],\mu_{\hat{a}_{k}}\geq\omega\mu_{1}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$ for $k\geq L$. The remaining work is to prove the upper bound of $\tau_{k}^{\text{ee}}$.
@@
From the algorithm design, we know the total pulling times in the exploration period up to phase k1k-1 is at most $T_{k-1}^{\text{ee}}$. By the Lemma B.5, we can validate the conditions of Lemma B.5 hold for $k\geq\left\{\kappa^{\text{ee}},\lceil\log_{2}\frac{\sum_{a=1}^{K}\frac{24}{\Delta_{0,a}^{2}}}{K}\rceil\right\}$. Thus, we have proved $\hat{a}_{k}=\textsf{None}$ for $k\geq\left\{\kappa^{\text{ee}},\lceil\log_{2}\frac{\sum_{a=1}^{K}\frac{24}{\Delta_{0,a}^{2}}}{K}\rceil\right\}$. The remaining work is to prove the upper bound of $N_{a}(\tau_{k}^{\text{ee}})$.
@@ C.3 Lower Bound for a Suboptimal Arm
where $Z_{1,s}$ is the realized KL divergence, $Z_{1,s}=\log\frac{\exp(-\frac{(X_{1,s}-\mu_{1})^{2}}{2})}{\exp(-\frac{(X_{1,s}-\tilde{\mu}_{1})^{2}}{2})}=\frac{(\mu_{1}-\tilde{\mu}_{1})(2X_{1,s}-\mu_{1}-\tilde{\mu}_{1})}{2}$, $X_{1,s}\sim N(\mu_{1},1)$. Or we can directly assume $Z_{1,s}\stackrel{{\scriptstyle i.i.d}}{{\sim}}N(\frac{(\mu_{1}-\tilde{\mu}_{1})^{2}}{2},(\mu_{1}-\tilde{\mu}_{1})^{2})$. Before moving on, we firstly prove a concentration result for the sum $\sum_{s=1}^{t}Z_{1,s}$. For any fixed number $N$, II, define $\xi_{N,I}=\left\{\max_{1\leq t\leq N}\sum_{s=1}^{t}\left(Z_{1,s}-\frac{(\mu_{1}-\tilde{\mu}_{1})^{2}}{2}\right)<I\right\}$. We can assert $\Pr_{Z_{1,s}\stackrel{{\scriptstyle i.i.d}}{{\sim}}N(\frac{(\mu_{1}-\tilde{\mu}_{1})^{2}}{2},(\mu_{1}-\tilde{\mu}_{1})^{2})}(\xi_{N,I})\geq 1-\exp\left(-\frac{I^{2}}{2N(\mu_{1}-\tilde{\mu}_{1})^{2}}\right)$, by the following application of maximal inequality (Theorem 3.10 in (Lattimore & Szepesvári, 2020)).
@@
$\stackrel{{\scriptstyle\lambda=\frac{I}{N(\mu_{1}-\tilde{\mu}_{1})^{2}}}}{{=}}$ $\Pr_{\tilde{Z}_{1,s}\stackrel{{\scriptstyle i.i.d}}{{\sim}}N(0,(\mu_{1}-\tilde{\mu}_{1})^{2})}\left(\max_{1\leq t\leq N}\prod_{s=1}^{t}\exp\left(\lambda\tilde{Z}_{1,s}\right)>\exp(\lambda I)\right)$
@@ D.1 Inequality about $x$ and $\log\log x$
$\stackrel{{\scriptstyle x>2\log x,\forall x>0}}{{>}}$ $0,$
@@ D.2 Probability Bound of Good Event
$\stackrel{{\scriptstyle\lambda=\frac{x}{u_{k}\sigma^{2}}}}{{=}}$ $\exp(-\frac{x^{2}}{2u_{k}\sigma^{2}}).$
@@ E.1 Settings of Numerical Experiments
We took arm number $K=10,20,30,40,50,100,150,200$ and the tolerance level $\delta=0.001,0.0001$. Fix $\mu_{0}=0.5,\Delta=0.15$, we set up 6 instances, by considering different number of arms whose mean rewards are above $\mu_{0}$. For an arm number $K$, we define (1) AllWorse, whose mean reward vector is $\mu_{1}=\mu_{2}=\cdots=\mu_{K}=0.25$; (2) Unique Qualified, whose mean reward vector is $\mu_{1}=\mu_{0}+\Delta,\mu_{2}=\mu_{3}=\cdots=\mu_{K}=\mu_{0}$; (3) One Quarter Qualified, whose mean reward vector is $\mu_{1}=\mu_{2}=\cdots=\mu_{\lfloor K/4\rfloor}=\mu_{0}+\Delta,\mu_{\lfloor K/4\rfloor+1}=\mu_{\lfloor K/4\rfloor+2}=\cdots=\mu_{K}=\mu_{0}$; (4) Half Good, whose mean reward vector is $\mu_{1}=\mu_{2}=\cdots=\mu_{\lfloor K/2\rfloor}=\mu_{0}+\Delta,\mu_{\lfloor K/2\rfloor+1}=\mu_{\lfloor K/2\rfloor+2}=\cdots=\mu_{K}=\mu_{0}$; (5) All Good, whose mean reward vector is $\mu_{1}=\mu_{2}=\cdots=\mu_{K}=\mu_{0}+\Delta$ (6) Linear, whose mean reward vector is $\mu_{i}=\mu_{0}-\Delta+\frac{2(i-1)\Delta}{K-1},1\leq i\leq K$. For instance “AllWorse”, the set of correct answer is $\{\textsf{None}\}$, while for the other instances which are positive, the answer set $i^{*}(\nu)$ contains at least one arm. In each experiment setting, we run 1000 independent trials and compute the empirical average of the stopping times.
@@ E.3 Correctness of APGAI algorithm
Instance Type $\setminus$ K $10$ 2020 $30$ $40$ $50$
@@
Instance Type $\setminus$ K $10$ 2020 $30$ $40$ $50$
@@
We tune some of the numeric experiment setting to see the non-stopping phenomenon more clearly. For APGAI, we set the forced stopping threshold as 50000K50000*K, where $K$ is the arm number. Once the total pulling times is no less than the forced stopping threshold, we mark this experiment as a failure, and take the forced stopping threshold as the total pulling times of APGAI. Table 2 records the number of failure for experiments $K=10,20,30,40,50$ and instance type “All Worse”, “Unique Qualified”, “One Quarter”, “Linear”. The tolerance level is $\delta=0.001$, with repeating times 1000. Except the negative instance “All Worse”, failure frequently occurs in some of positive instances. In the “Unique Qualified” group, at least 35% of experiments end up with “failure”. While in other groups, failure also occurs. In group “One Quarter”, the number of failures decreases as arm number $K$ increases, suggesting APGAI’s good performance relies more on the number of qualified arms, instead of the fraction of qualified arms.